OpenStudy (anonymous):

The sum of an integer and its reciprocal is 65/8. Find the integer

OpenStudy (anonymous):

\[x + \frac{ 1 }{ x } = \frac{ 65 }{8 }\] multiply by x and its the same idea

OpenStudy (anonymous):

how do i multiply x by 65/8

OpenStudy (mathstudent55):

Multiply both sides by 8x to get rid of the fractions.

OpenStudy (anonymous):

so is it 8,1/8

OpenStudy (mathstudent55):

\(x + \dfrac{1}{x} = \dfrac{65}{8} \) Multiply both sides by 8x: \((8x)(x + \dfrac{1}{x}) = (8x)(\dfrac{65}{8}) \) \((8x)(x) + (8x)(\dfrac{1}{x}) = (8x)(\dfrac{65}{8}) \) \((8x)(x) + (8\cancel{x})(\dfrac{1}{\cancel{x}}) = (\cancel{8}x)(\dfrac{65}{\cancel{8}}) \) \(8x^2 + 8 = 65x \) \(8x^2 - 65x + 8 = 0 \) Now either try factoring or use the quadratic formula. The answer is 8 or 1/8. You are correct.

OpenStudy (anonymous):

thanks mathstudent your the man

OpenStudy (mathstudent55):

You're welcome.